OpenStudy (anonymous):

Find the volume of the solid formed by rotating the region inside the first quadrant enclosed by y=x^3 and y=9x about the x-axis.

OpenStudy (zarkon):

first solve \[x^3=9x\]

OpenStudy (anonymous):

x(x^2-9)=0

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

so x=?

OpenStudy (anonymous):

x=0 at 0 and 3?

OpenStudy (zarkon):

0, -3 ,3 ....but we can ignore the -3 since it is not in the first quadrant.

OpenStudy (zarkon):

so our integral is going to go from x=0 to x=3

OpenStudy (zarkon):

between 0 and 3 , which function is on top?

OpenStudy (anonymous):

x^3 should be and then subtract 9x?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

pick any number between 0 and 3...like 1 and plug into each function 1^3=1 9*1=9 9>1

OpenStudy (zarkon):

so 9x is on top of x^3 between 0 and 3

OpenStudy (zarkon):

now we can use the formula \[V=\int_{a}^{b}\pi f(x)^2dx\]

OpenStudy (anonymous):

so from 0,3 pi(9x)^2- pi(x^3)^2= VOLUME?

OpenStudy (zarkon):

\[\int\limits_{0}^{3}(\pi (9x)^2-\pi(x^3)^2)dx\] yes

OpenStudy (anonymous):

so (pi(9x^3/3)-pi(x^7/7))dx

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[\int\limits_{0}^{3}(\pi (9x)^2-\pi(x^3)^2)dx\] \[=\pi\int\limits_{0}^{3}(81x^2-x^6)dx\]

OpenStudy (anonymous):

the volume is\[2\times \int\limits_{0}^{3}[\pi(9x)^2-\pi(x^3)^2]dx\]

OpenStudy (zarkon):

we are working in the first quadrant

OpenStudy (anonymous):

so i won't have to take the antiderivative?

OpenStudy (zarkon):

you will....you forgot to square your 9

OpenStudy (zarkon):

\[=\pi\left[81\frac{x^3}{3}-\frac{x^7}{7}\right]_{0}^{3}\]

OpenStudy (zarkon):

\[=\pi\left[81\frac{3^3}{3}-\frac{3^7}{7}\right]\]

OpenStudy (anonymous):

and 2×∫ 0 3 [π(9x) 2 −π(x 3 ) 2 ]dx =5832pi/7

OpenStudy (zarkon):

you don't need the times 2.

OpenStudy (anonymous):

416.5714pi

OpenStudy (zarkon):

\[=\frac{2916\pi}{7}\]

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thank you so much

OpenStudy (zarkon):

no problem

OpenStudy (anonymous):

?the graph is adove x-axis and below x-axis,it is symetric,so i have to times 2

OpenStudy (zarkon):

no you don't

OpenStudy (zarkon):

all we need it the area of a circle \[\pi r^2\] if f(x) is the height of the function then we have \[\pi f(x)^2\] we can then multiply by the thickness of our disks and add them all up V=\[\int\limits_{a}^{b}\pi (f(x))^2dx\] no 2 times in this formula.

OpenStudy (anonymous):